# May the Feedback Be with You! Unlocking the Power of Feedback-Driven Deep Learning Framework Fuzzing via LLMs

### v2

@@ 5.4. Ablation Studies
Similar to the observations of prior techniques (FuzzGPT (Deng et al., 2024b), WhiteFox (Yang et al., 2024)), we notice that the PyTorch community is more active than the TensorFlow community as PyTorch developers confirm and fix issues faster. Therefore, we run our ablation experiments on PyTorch over default fuzzing configuration. Table 4 shows ablation comparison results.
@@ 6.2. DL Framework Testing
One of the pioneering works is CRADLE (Pham et al., 2019), which detects inconsistencies by running pre-built DL models across multiple low-level backends of Keras (Keras, 2020). To enhance the diversity of models, LEMON (Wang et al., 2020) and AUDEE (Guo et al., 2020) build upon CRADLE by applying pre-defined mutation rules to seed models and inputs, generating a broader range of test cases. NNSmith (Liu et al., 2023a) models each DL operator with input constraints and shape inference, using SMT solving to generate valid models. Subsequently, NeuRI (Liu et al., 2023b) is proposed to synthesize DL models via rule inference automatically. Unfortunately, the above traditional fuzzers need manual design for mutation rules, operator specifications, or synthesis rules, which need domain-specific expert knowledge. With the success of LLMs in software testing (Lemieux et al., 2023; Feng and Chen, 2024), researchers have attempted to apply LLMs to fuzzing DL frameworks. TitanFuzz (Deng et al., 2023) is the first work to directly leverage LLMs to generate API call sequences and infill these APIs. Then FuzzGPT (Deng et al., 2024b) is proposed to generate edge-case code via in-context learning and fine-tuning. WhiteFox (Yang et al., 2024) proposes a white-box compiler fuzzing technique aided by LLMs, targeting to generate specific inputs to trigger the optimization features in compilers. However, all of them struggle to generate diverse and valid DL models, and FUEL is proposed to fill this gap by adopting a heuristic search for operator selection and a program self-repair strategy by mining the potential of feedback.
@@ 7. Conclusion
Acknowledgment
We thank Wei Cheng, Yanzhou Mu, and Zhiyuan Li for their insightful feedback on the manuscript.

### v3

@@ 5.4. Ablation Studies
Similar to the observations of prior studies (FuzzGPT (Deng et al., 2024b), WhiteFox (Yang et al., 2024)), we notice that the PyTorch community is more active than the TensorFlow community as PyTorch developers confirm and fix issues faster. Therefore, we run our ablation experiments on PyTorch with the default fuzzing configuration. Table 4 shows ablation comparison results.
@@ 6.2. DL Framework Testing
One of the pioneering works is CRADLE (Pham et al., 2019), which detects inconsistencies by running pre-built DL models across multiple low-level backends of Keras (Keras, 2020). To enhance the diversity of models, LEMON (Wang et al., 2020) and AUDEE (Guo et al., 2020) build upon CRADLE by applying pre-defined mutation rules to seed models and inputs, generating a broader range of test cases. NNSmith (Liu et al., 2023a) models each DL operator with input constraints and shape inference, using SMT solving to generate valid models. Subsequently, NeuRI (Liu et al., 2023b) is proposed to synthesize DL models via rule inference automatically. Unfortunately, the above traditional fuzzers need manual design for mutation rules, operator specifications, or synthesis rules, which need domain-specific expert knowledge. With the success of LLMs in software testing (Lemieux et al., 2023; Feng and Chen, 2024), researchers have attempted to apply LLMs to fuzzing DL frameworks. TitanFuzz (Deng et al., 2023) is the first work to directly leverage LLMs to generate API call sequences. Then FuzzGPT (Deng et al., 2024b) is proposed to generate edge-case code via in-context learning and fine-tuning. WhiteFox (Yang et al., 2024) proposes a white-box compiler fuzzing technique aided by LLMs, targeting to generate specific inputs to trigger the optimization features in compilers. However, all of them struggle to generate diverse and valid DL models, and FUEL is proposed to fill this gap by adopting a heuristic search for operator selection and a program self-repair strategy by mining the potential of feedback.